1.  At 00C, a rod of aluminum and a rod of copper have the same length, 1.25 meters. To what common temperature must they be heated to differ in length by 0.0750 millimeter?



1
Expert's answer
2021-02-07T19:22:37-0500

By the definition of thermal linear expansion, we have:


"\dfrac{L_{Al}-L_{Al,0}}{L_{Al,0}}=\alpha_{Al}(t_{comm}-0^{\circ}C),""\dfrac{L_{Cu}-L_{Cu,0}}{L_{Cu,0}}=\alpha_{Cu}(t_{comm}-0^{\circ}C)."

Then, taking into account that "L_{Al,0}=L_{Cu,0}=L_0" we can rewrite our equations as follows:


"L_{Al}-L_0=\alpha_{Al}L_0t_{comm},""L_{Cu}-L_0=\alpha_{Cu}L_0t_{comm}."

Subtracting second equation from the first one, we get:


"L_{Al}-L_{Cu}=(\alpha_{Al}-\alpha_{Cu})L_0t_{comm},""t_{comm}=\dfrac{L_{Al}-L_{Cu}}{(\alpha_{Al}-\alpha_{Cu})L_0},""t_{comm}=\dfrac{7.5\cdot10^{-5}\ m}{(24\cdot10^{-6}\ \dfrac{1}{^{\circ}C}-17\cdot10^{-6}\ \dfrac{1}{^{\circ}C})\cdot1.25\ m}=8.57^{\circ}C."